# Are you sure? Measuring models bias in content moderation through uncertainty

### v1

@@ 5 STUDY 1: Models Uncertainty towards Socio-Demographic Groups
IMSyPP 0.41 (0.03) +58e-3 -17e-3 -4e-2 +1e-2 0.33 (0.01) -16e-3 +1e-3 -7e-4 +6e-3
HateBert 0.51 (0.04) -31e-4 -23e-4 55e-4 55e-4 0.49 (0.01) -2e-3 +1e-3 +1e-3 +3e-4
Dynabench 0.29 (0.02) -5e-3 +5e-3 -26e-3 +35e-3 0.34 (0.01) -4e-4 +13e-3 +8e-3 +5e-3
Twitter-Roberta-Base 0.31 (0.03) +1e-3 +8e-3 -19e-3 +27e-3 0.37 (0.01) -3e-3 -1e-2 +4e-3 +3e-3
Refugees 0.57 (0.02) -7e-3 +6e-3 +29e-3 -9e-3 0.55 (0.02) +4e-3 +2e-3 -38e-4 -24e-4
DistilRoberta 0.44 (0.03) -11e-3 +6e-3 -19e-3 +1e-2 0.44 (0.01) -1e-2 +9e-3 +6e-4 -4e-4
Pysentimiento 0.35 (0.02) 8e-3 2e-3 -58e-4 +26e-3 0.33 (0.02) -19e-3 +13e-3 +13e-3 +9e-3
MuRIL 0.36 (0.03) 10e-3 +12e-6 -55e-4 12e-3 0.31 (0.01) -17e-3 +12e-3 +6e-4 +5e-4
Olmo-7B 0.48 (0.04) -68e-4 +35e-5 +51e-3 -12e-3 0.55 (0.03) 12e-3 -44e-4 -74e-4 -21e-4
Bloom-7B 0.48 (0.03) -30e-4 -14e-4 +19e-3 +16e-3 0.49 (0.02) +22e-4 -65e-4 +6e-4 +11e-4
Mistral-7B 0.49 (0.03) -23e-4 +31e-4 -21e-4 +53e-4 0.50 (0.03) -67e-4 +75e-4 +24e-4 -40e-4
Table 1: Delta $F_{1}$ score for SBIC and CREhate obtained with each model against the total list of disaggregated labels and the $F_{1}$ score against labels of a specific group: white men ($w.m.$); white women ($w.f.$); non-white men ($\neg w.m.$), non-white women ($\neg w.f.$).
@@
Table 2: The Uncertainty Divergence divergence between the distribution of non-conformity scores assigned by models to all the individual annotations against the distributions of annotations grouped by the intersection of gender and ethnicity. The highest is the score the highest is the divergence between the non-conformity of all the individual annotations and socio-demographic groups.
@@ Appendix C Cohen’s Kappa between annotators pairs
Figure 6: Pairwise Cohen’s Kappa agreement between annotator classes per CREHate.

### v2

@@ 5 STUDY 1: Models Uncertainty towards Socio-Demographic Groups
IMSyPP 0.41 +58e-3 -17e-3 -4e-2 +1e-2 0.33 -16e-3 +1e-3 -7e-4 +6e-3
HateBert 0.51 -31e-4 -23e-4 55e-4 55e-4 0.49 -2e-3 +1e-3 +1e-3 +3e-4
Dynabench 0.29 -5e-3 +5e-3 -26e-3 +35e-3 0.34 -4e-4 +13e-3 +8e-3 +5e-3
Twitter-Roberta-Base 0.31 +1e-3 +8e-3 -19e-3 +27e-3 0.37 -3e-3 -1e-2 +4e-3 +3e-3
Refugees 0.57 -7e-3 +6e-3 +29e-3 -9e-3 0.55 +4e-3 +2e-3 -38e-4 -24e-4
DistilRoberta 0.44 -11e-3 +6e-3 -19e-3 +1e-2 0.44 -1e-2 +9e-3 +6e-4 -4e-4
Pysentimiento 0.35 8e-3 2e-3 -58e-4 +26e-3 0.33 -19e-3 +13e-3 +13e-3 +9e-3
MuRIL 0.36 10e-3 +12e-6 -55e-4 12e-3 0.31 -17e-3 +12e-3 +6e-4 +5e-4
Olmo-7B 0.48 -68e-4 +35e-5 +51e-3 -12e-3 0.55 12e-3 -44e-4 -74e-4 -21e-4
Bloom-7B 0.48 -30e-4 -14e-4 +19e-3 +16e-3 0.49 +22e-4 -65e-4 +6e-4 +11e-4
Mistral-7B 0.49 -23e-4 +31e-4 -21e-4 +53e-4 0.50 -67e-4 +75e-4 +24e-4 -40e-4
Table 1: Delta $F_{1}$ score for SBIC and CREhate obtained with each model against the total list of disaggregated labels and the $F_{1}$ score against labels of a specific group: white men ($w.m.$); white women ($w.f.$); non-white men ($\neg w.m.$), non-white women ($\neg w.f.$). Models that encode optimally the perspective of a group are highlighted in green. Models that predict worse on the group than on majority voting are highlighted in red.
@@
Table 2: The Uncertainty Divergence between the distribution of non-conformity scores assigned by models to all the individual annotations against the distributions of annotations grouped by the intersection of gender and ethnicity. The higher the score (red), the higher the divergence between the non-conformity of all the individual annotations and socio-demographic groups. The lower the score (green), the lower the divergence.
@@ Appendix C Cohen’s Kappa between annotators pairs
Figure 6: Pairwise Cohen’s Kappa agreement between annotator classes per CREAHate.